(15’) How many ml of 2.00 M HNO3 would be required to neutralize 12.5 ml of 0.0800 M NH3? What is the pH of the resulting solution? (KaNH4=5.6x10-10)


1
Expert's answer
2020-11-11T10:36:26-0500

NH3 + HNO3 = NH4NO3

CaVa/CbVb = na/nb

2×Va/ 0.08× 12.5 = 1/1

Va= 0.08×12.5/2

Va= 0.5mL


Mole of HNO3 and NH3= mole of NH4NO3

mole of HNO3 = 2×0.5×10-3 = 1×10-3 mol

Concentration of NH4NO3 = 1×10-3mol/[(0.5+12.5) × 10-3 L]

= 0.0769M


PH of solution= 1/2 Pka - 1/2 log c

= 1/2(-log 5.6×10-10 )- 1/2 log 0.0769

=5.18